# Data Trade and Consumer Privacy

### v2

@@ 4.4 Welfare and Privacy Protection
Table 4 summarizes social welfare with and without data trade market and Table 5 shows changes in welfare. Banning data market decreases social welfare when $\mu_{0}>\frac{v_{L}}{v_{H}}$. Intuitively, without data trade, it becomes harder for the producer to price discriminate consumers and extract the most surplus. Specifically, the data trade brings two effects: surplus extraction and market inclusion. The surplus extraction effect does not matter for social welfare as it is a reallocation of surplus between the producer and consumers. The market inclusion effect of data trade is $\left(1-\mu_{0}\right)v_{L}\min\left\{\Delta v,1\right\}$ when the proportion of the type-L consumers is large. This positive effect will disappear if laws ban the data market.
Proposition 4.
Banning data trade decreases social welfare if the proportion of the type-H consumers in the market is large $\mu_{0}>v_{L}/v_{H}$.
Table 4: Welfare Summary
Prior Mismatch With Data Market Without Data Market
$\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}v_{H}$ $V+\mu_{0}v_{H}$
$\delta^{2}<\Delta v$ $V+\left(1-\mu_{0}\right)v_{L}\Delta v+\mu_{0}v_{H}$ $V+\mu_{0}v_{H}$
$\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$ $V+v_{L}+\mu_{0}\Delta v$ $V+v_{L}+\mu_{0}\Delta v$
$\delta^{2}<\Delta v$ $V+v_{L}+\mu_{0}\Delta v$ $V+v_{L}+\mu_{0}\Delta v$
Table 5: Change in Welfare Led by Ban
Prior Mismatch Welfare Change
$\mu_{0}>\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$downarrow
$\delta^{2}<\Delta v$downarrow
$\mu_{0}<\frac{v_{L}}{v_{H}}$ $\delta^{2}>\Delta v$-
$\delta^{2}<\Delta v$-
5 General Service
The previous baseline analysis is under the binary service. As the mismatch level in service changes consumers’ incentives, this section investigates a general service space $\mathcal{X}=\mathbb{R}$ which allows flexible mismatch level. The analysis shows that the main insights hold with a general service. The general service generates a different cutoff for selling full information to be optimal.
The results rely on the market prior similar to the baseline. When $\mu_{0}<v_{L}/v_{H}$, there is no distortion in the intermediary optimal service provision even under a general service space. The results on service fee, data trade, and data fee are then as Proposition 2: selling no information can be optimal. We relegate them and their proofs to Appendix A.8 of Proposition 7.
We summarize the results for $\mu_{0}>v_{L}/v_{H}$ as follows.
Proposition 5.
Suppose the service space $\mathcal{X}=\mathbb{R}$ and $\mu_{0}>v_{L}/v_{H}$. In the optimal mechanisms, all the consumers surpluses are extracted. For $\delta^{2}\geq\Delta v$, there is no distortion in service provisions $x_{L}=L$, $x_{H}=H$, and the service fees, data trade are as the part 1 of Proposition 1.
For $\delta^{2}<\Delta v$, there is a downward distortion in service provided to type-$L$ consumers, but there is no service distortion for type-$H$ consumers $x_{H}=H$. The following hold:
@@ 5 General Service
    $\pi(l|L)=1,\text{ }\pi(l|H)\in[0,C]$
    the data trade fee is
    $T=-\mu_{0}\pi(l|H)\Delta v+(1-\mu_{0})v_{L}$
    and the intermediary’s total revenue is
    $V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}$
    In this case social welfare is
    $V-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}+(1-\mu_{0})v_{L}+\mu_{0}v_{H}$
@@
    If $\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{L}+v_{H}}$, then the service for type-$L$ is $x_{L}=L-\frac{\delta}{\Delta v}v_{L}$, the service fees are $f_{L}=V-(\frac{\delta}{\Delta v}v_{L})^{2}$, $f_{H}=V+\pi(l|H)\Delta v$; and the data trade is
    $\pi(l|L)=\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v},\text{ }\pi(l|H)\in\Big{[}0,C\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}\Big{]}$
    the data trade fee is
    $T=-\mu_{0}\pi(l|H)\Delta v+(1-\mu_{0})v_{L}\frac{\delta^{2}}{(\Delta v)^{2}}(v_{L}+v_{H})$
    and the intermediary’s total revenue is
    $V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}.$
    In this case social welfare is
    $V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}+\mu_{0}v_{H}.$
When $\mu_{0}>v_{L}/v_{H}$ and the mismatch cost is large enough, i.e., $\delta^{2}>\Delta v$, there is no distortion in consumers’ service even with a general service space. This is because the mismatch cost in service is so high that consumers are willing to report the private information without any further distortion in service level. The intermediary sets the highest $\pi\left(l|L\right)=1$ and does not need to distort any service level to maintain the type-H’s incentive. The results on service fees and data trade are as the part 1 of Proposition 1.
However, when $\mu_{0}>v_{L}/v_{H}$ and the mismatch cost is small, $\delta^{2}<\Delta v$, the intermediary needs to downward distort the type-L’s service. In this case, the penalty of the wrong service cannot provide enough incentive. The intermediary will manipulate the type-H consumers’ incentive via other two instruments: the probability of surplus in the product and service provision. Although distorting the service will reduce the surplus extracted from the type-Ls in the service market, the intermediary can extract more surplus from the type-H consumers in the product market. When the benefit of more surplus from type-H consumers outweighs the cost of reduced surplus from type-L consumers, the intermediary would prefer to distort the service to the type-L consumers. This happens when the population of type-H consumers is large enough $\mu_{0}>v_{L}/v_{H}$.
In sum, the results on data trade are similar to the baseline analysis, but with a different threshold for selling full information to be optimal. Under a general service space selling full information can be optimal when the mismatch cost is sufficiently large, i.e., $\delta^{2}>(\Delta v)^{2}/(v_{L}+v_{H})$ by the part 2 of Proposition 5. The cutoff $(\Delta v)^{2}/(v_{L}+v_{H})$ is smaller than the one, $\Delta v$ in baseline results under binary service.
As both service distortion and price discrimination affect consumers’ incentives, the general service setup allows us to discuss their relations in the optimal mechanisms. The below follows directly from propositions 5 and 7.
Corollary 4 (Service distortion and price discrimination).
In the intermediary optimal mechanisms, the service distortion and no-price discrimination are exclusive.
Given a service distortion, we have $\mu_{0}>v_{L}/v_{H}$ and $\delta^{2}<\Delta v$; otherwise, there is no distortion by Proposition 5 and 7. From $\mu_{0}>v_{L}/v_{H}$ and $\delta^{2}<\Delta v$, it is impossible that $\pi(l|L)=\pi(l|H)$, which implies no-discrimination is not possible.
Instead, given no price discrimination, it has to be $\mu_{0}<v_{L}/v_{H}$ by contradiction: Suppose $\mu_{0}>v_{L}/v_{H}$, then it is impossible for $\pi(l|H)=\pi(l|L)$. Hence, $\mu_{0}<v_{L}/v_{H}$ implies that there is no service distortion from Proposition 7.
Regulation.
Similarly, we discuss the welfare implications of banning data trade within this more general setup. The main welfare implications are same with those in the binary service and are summarized into the following proposition.
Proposition 6.
Suppose a general service space. Banning data trade will decrease social welfare for $\mu_{0}>v_{L}/v_{H}$, and will not affect social welfare for $\mu_{0}<v_{L}/v_{H}$.
@@ 6 Discussion
In the stylized model, there is one data buyer, i.e., the producer. When there are two producers and both of them buy data from the intermediary, the pricing competition in the product market between two producers might compete away the producer’s profit and hence prevent the intermediary from a beneficial data trade. In this sense, the intermediary would prefer an exclusive data trade, selling data only to one producer, which goes back to the baseline framework of one data buyer.
@@
Recall that our work focuses on the cross-market incentive instead of the inter-temporal one. If one interprets our results from the inter-temporal viewpoint, there can be a commitment problem of the intermediary under some conditions. Take the result in Corollary 3 as an example, when $\delta^{2}<\Delta v$, the data trade ex ante committed by the optimal mechanism can be inconsistent with what will happen after consumers report the true types in the service market. The optimal mechanism will never sell full information when $\delta^{2}<\Delta v$ while the intermediary has incentives to sell all the information to get the maximum data trade gain. [^24] The intermediary’s commitment is not a problem when $\delta^{2}$ is large enough. The full information data trade is optimal if only if $\delta^{2}$ is sufficiently large. Given a full information data trade in the optimality, there will not be the time-inconsistent behaviors of the intermediary after collecting the data. Bester and Strausz (2001), Doval and Skreta (2022) and Doval and Skreta (Forthcoming) study the mechanism with the principal’s limited commitment.
@@ 7 Conclusion
This paper develops an integrated framework of the intermediary’s data sourcing from consumers and data trade to a third party: Consumers benefit from revealing information about their taste to the intermediary for better-matched service. A third party buys data from the intermediary so that it can extract more surplus from consumers in the product market. We characterize the optimal mechanisms for collecting and selling data. Due to the cross-market role of data, the equilibrium data trade depends on the consumers’ horizontal differentiation level in the data-sourcing service market and the vertical one in the product market.
In the intermediary optimal mechanism, the intermediary sells no information or partial information to preserve profits in the service market when consumers’ mismatch cost from a wrong service is low. The low mismatch cost itself cannot sufficiently incentivize consumers to reveal information. The intermediary has to compensate consumers more (by decreasing service fees) to sell full information. However, the loss in service profit outweighs the net gain from selling full information rather than partial or no information. Hence, the intermediary would prefer to sell partial or no information, which suggests that the market can protect consumer privacy without regulations. In contrast, the intermediary can sell full information to data buyers when consumers’ mismatch cost in service is high. In this case, consumers’ privacy protection is at the lowest level.
Using the framework to evaluate privacy protection, we find that a ban on data trade may reduce social welfare because it makes it harder to price discriminate in the product market. The finding suggests an alternative policy to regulate the data market, such as a lump-sum tax. In this way, there will not be welfare loss, and consumers and the producer can share some surplus with the intermediary via government transfers.
Appendix A Appendix
A.1 Proof of Lemma 1

### v3

@@ 4.3 Regulation of Data Sharing
Social welfare is the sum of the intermediary’s profit, producer’s profit, and consumer surplus in Table 2. Hence, banning the data market decreases social welfare when $\mu_{0}>v_{L}/v_{H}$. [^24] Acemoglu et al. (2022) find that banning the data market improves welfare as the over-sharing of data in the market induced by data externality disappears. The welfare loss comes from the firm’s over-collection of consumer data in Bergemann et al. (2022) and Choi et al. (2019). Intuitively, without data trade, it becomes harder for the producer to price discriminate consumers and extract the most surplus. Specifically, the data trade brings two effects: surplus extraction and market inclusion. The surplus extraction effect does not matter for social welfare as it is a redistribution of surplus between the producer and consumers. But the positive market inclusion effect will disappear if the data trade is banned.
Proposition 4 (Regulation of data sharing).
Banning data trade will decrease social welfare for $\mu_{0}>v_{L}/v_{H}$, and will not affect social welfare for $\mu_{0}<v_{L}/v_{H}$.
The result suggests that the privacy laws should be careful about the levels of protection. Depending on the market environment, the highest privacy protection is sometimes optimal (e.g., $\mu_{0}<v_{L}/v_{H}$), while trading full data is sometimes optimal (e.g., $\delta^{2}\geq\Delta v$).
5 Intermediary’s Bargaining Power
One assumption in the baseline analysis of intermediary optimal design is that the intermediary has the full bargaining power in the data market and hence extracts all the data trade surplus. However, transferring data from consumers to the intermediary does not necessarily fully transfer the bargaining power to the latter. This section studies the partial bargaining power of the intermediary and its regulatory implications.
In a stochastic ultimatum bargaining between the intermediary and the producer in the data market, the intermediary proposes a take-it-or-leave-it offer with probability $\beta\in(0,1)$, which measures the intermediary’s relative bargaining power. When the intermediary extracts a proportion $\beta$ of the data trade surplus, its total payoff becomes
$f_{H}\mu_{0}+f_{L}(1-\mu_{0})+\beta[(1-\mu_{0})\pi(l|L)v_{L}-\mu_{0}\pi(l|H)(v_{H}-v_{L})+(\mu_{0}v_{H}-v_{L})\mathbb{I}_{\mu_{0}<v_{L}/v_{H}}],$
while the constraints of the problem are unchanged. Solving the problem yields results about the intermediary’s optimality with partial bargaining power.
Proposition 5 (Optimal design—partial bargaining power).
Suppose a partial bargaining power of the intermediary. In the intermediary optimal mechanism:
@@ 5 Intermediary’s Bargaining Power
    For $\mu_{0}>v_{L}/v_{H}$, there is no distortion of the service for the high type $x_{H}=H$ and
    1. (a)
      if $\frac{\delta^{2}}{\Delta v}\geq 1$, then the data trade $\pi(l|L)=1$, $\pi(l|H)=C$ and the service is $x_{L}=L$;
    2. (b)
      if $\frac{\Delta v}{v_{H}+v_{L}}<\frac{\delta^{2}}{\Delta v}<1$, then the data trade is $\pi(l|L)=1$, $\pi(l|H)=C$ and the service is $x_{L}=L-\frac{\Delta v-\delta^{2}}{2\delta}$;
    3. (c)
      if $\frac{\delta^{2}}{\Delta v}\leq\frac{\Delta v}{v_{H}+v_{L}}$, then the data trade is $\pi(l|L)=\frac{\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}$, $\pi(l|H)=\frac{C\delta^{2}}{\Delta v}\frac{v_{H}+v_{L}}{\Delta v}$ and the service is $x_{L}=L-\frac{\delta v_{L}}{\Delta v}$.
@@
    For $\mu_{0}<v_{L}/v_{H}$, the optimal data trade is $\pi\left(l|L\right)=\pi\left(l|H\right)=1$, and services are $x_{L}=L$, $x_{H}=H$.
Comparing the intermediary’s optimality under the partial bargaining power with that under the full bargaining power in the propositions 2 and 3, two main results follow. First, the partial bargaining power can decrease the amount of traded data. The optimal $\pi(l|H)$ under the partial bargaining power becomes unique from multiple solutions. For example, in the case of $\mu_{0}>v_{L}/v_{H}$, fixing the optimal $\pi(l|L)$, the data trade changes from $\{\pi(l|L),\pi(l|H)\in[0,C\pi(l|L)]\}$ to $\{\pi(l|L),\pi(l|H)=C\pi(l|L)\}$ with $C\in(0,1)$. [^25] When $\mu_{0}>\frac{v_{L}}{v_{H}}$, in the optimal mechanism (16) is binding, $\pi(l|H)=C\pi(l|L)$, which together with other constraints pins down the unique solution. Hence, the partial bargaining power drives down the amount of information disclosed by the intermediary’s optimal mechanism.
Second, the partial bargaining power does not affect the intermediary’s equilibrium payoff, although it shuts off the multiple solutions of data trade. [^26] In the optimal mechanism, the intermediary’s equilibrium payoff is $V$ for $\mu_{0}<\frac{v_{L}}{v_{H}}$, and it is $\begin{cases}V+(1-\mu_{0})v_{L}&\mbox{ if }\frac{\delta^{2}}{\Delta v}\geq 1\\ V+(1-\mu_{0})v_{L}-(1-\mu_{0})\left(\frac{\Delta v-\delta^{2}}{2\delta}\right)^{2}&\mbox{ if }\frac{\Delta v}{v_{H}+v_{L}}\leq\frac{\delta^{2}}{\Delta v}<1\\ V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}&\mbox{ if }\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{H}+v_{L}}\end{cases}$ for $\mu_{0}>\frac{v_{L}}{v_{H}}$, which is exactly the intermediary’s equilibrium payoff in the setup of $\beta=1$. In the case of $\mu_{0}>v_{L}/v_{H}$, the intermediary’s total payoff can be written as $V-(x_{H}-H)^{2}+\pi(l|H)\Delta v]\mu_{0}+[V-(x_{L}-L)^{2}](1-\mu_{0})+\beta[(1-\mu_{0})\pi(l|L)v_{L}-\mu_{0}\pi(l|H)\Delta v]$ where in the optimal mechanism the data revenue is zero because
$(1-\mu_{0})\pi(l|L)v_{L}-\mu_{0}\pi(l|H)\Delta v$ $=$ $(1-\mu_{0})\pi(l|L)v_{L}-\mu_{0}C\pi(l|L)\Delta v$
$=$ $(1-\mu_{0})\pi(l|L)v_{L}-(1-\mu_{0})\pi(l|L)v_{L}$
$=$ $0$
where the first equality follows the binding obedience constraint $OB_{L}$ in (16): $\pi(l|H)=C\pi(l|L)$ and the second follows the expression of $C$ in 1. Hence, the bargaining power parameter doesn’t matter for the equilibrium payoff when $\mu_{0}>v_{L}/v_{H}$.
In the case of $\mu_{0}<v_{L}/v_{H}$, the intermediary’s total payoff reads as
$\underbrace{[V-(x_{H}-H)^{2}+\pi(l|H)\Delta v-\Delta v]}_{\text{service fee for type-H consumer}}\mu_{0}+[V-(x_{L}-L)^{2}](1-\mu_{0})$
$\hskip 144.54pt+\beta[(1-\mu_{0})\pi(l|L)v_{L}-\mu_{0}\pi\left(l|H\right)\Delta v-(v_{L}-\mu_{0}v_{H})]$
There is a tradeoff of increasing the probability of charging a low price for the high type, $\pi(l|H)$. As $\pi(l|H)$ increases, on one hand, the service revenue increases by $\mu_{0}\Delta v$ because consumers are more likely to reveal personal information. On the other hand, the data revenue decreases by $\beta\mu_{0}\Delta v$ because the producer charges the type-H consumers a low price with a higher probability. The first force outweighs the second given the partial bargaining power $\beta<1$. So the intermediary will increase $\pi(l|H)$ up to one. [^27] Recall that when $\beta=1$, the two forces cancel each other out and hence $\pi(l|H)$ is free to have multiple solutions. Moreover, the intermediary sets $\pi(l|L)=1$ to maximize the data revenue. Hence, the optimal data revenue is zero because $(1-\mu_{0})\pi(l|L)v_{L}-\mu_{0}\pi\left(l|H\right)\Delta v-(v_{L}-\mu_{0}v_{H})=0$. The bargaining power $\beta<1$ does not influence the intermediary’s equilibrium total payoff, which is $V$.
Regulation of Power.
Recall that the consumer surplus is $\max\{0,\mu_{0}\Delta v\mathbb{I}_{\mu_{0}v_{H}<v_{L}}\}$ and the producer surplus is $\max\{v_{L},\mu_{0}v_{H}\}$. So the partial bargaining power of the intermediary does not affect the consumer surplus, producer profit, nor the total welfare, compared to the full bargaining power. In sum, the introduction of the partial bargaining power yields the following insights about the regulation of intermediary’s power in the data market.
Proposition 6 (Regulation of intermediary’s power).
Reducing the intermediary’s bargaining power in the data market can decrease the amount of information traded by the intermediary, without hurting social welfare.
@@ 6 Discussion
Other Off-path Beliefs.
The equilibrium selection criterion restricts the off-equilibrium-path belief to be $\mu_{0}$, i.e., the producer does not update from the consumer’s non-participation. This is a restriction of “no signaling what you don’t know” as in Fudenberg and Tirole (1991). Alternatively, for example, the producer may treat the consumers as high type if they reject the intermediary’s contract, i.e., the off-path belief is $1$. Then the producer sets the product price $p_{0}=v_{H}$ and $R_{0}=v_{H}$, the type-H consumers’ outside option is 0. Recall in the analysis of the off-path belief being $\mu_{0}$, when $\mu_{0}>v_{L}/v_{H}$, the producer sets product price at $v_{H}$ and the revenue is $\mu_{0}v_{H}$, the type-H consumers’ outside option is 0. So this particular off-path belief changes the right-hand side of the producer’s participation constraint (9) from $\mu_{0}v_{H}$ to $v_{H}$, and the other constraints do not change. Hence, the results on data trade in the case of $\mu_{0}>v_{L}/v_{H}$ won’t change, while the data fee will change. Off-path belief restrictions are also seen in related models like Lizzeri (1999).
@@
In the stylized model, there is one data buyer, i.e., the producer. When there are two producers and both of them buy data from the intermediary, the pricing competition in the product market between the two producers might compete away the producer’s profit and hence prevent the intermediary from a beneficial data trade. In this sense, the intermediary would prefer an exclusive data trade, selling data only to one producer, which goes back to the baseline framework of one data buyer. Bimpikis et al. (2019) and Bonatti et al. (2024) study information sales to buyers who compete in a downstream market. Armstrong and Zhou (2022) study the intermediary designed information based on which the information users compete in the product price.
@@
Recall that our work focuses on the cross-market incentive instead of the inter-temporal one. If one interprets our results from the inter-temporal viewpoint, there can be a commitment problem of the intermediary under some conditions. Take the result of $\mu_{0}<v_{L}/v_{H}$ in 2 as an example, when $\delta^{2}<\Delta v$, the data trade ex ante committed by the optimal mechanism can be inconsistent with what will happen after consumers report the true types in the service market. The optimal mechanism will never sell full information when $\delta^{2}<\Delta v$, but the intermediary has incentives to sell all the information to get the maximum data trade gain. [^28] In the case of $\mu_{0}<v_{L}/v_{H}$, the intermediary’s commitment is not a problem when $\delta^{2}\geq\Delta v$: the full information data trade is optimal if $\delta^{2}\geq\Delta v$ and hence there will be no time-inconsistent behaviors of the intermediary after collecting the data. Bester and Strausz (2001), Doval and Skreta (2022) and Doval and Skreta (2025) study the mechanism with the principal’s limited commitment.
Transparency of Contract.
The paper assumes consumers observe everything about the data contract offered to the data-using producer. It is not an unreasonable assumption because transparency is one of the key aspects of data regulations, i.e., businesses must be transparent about their data collection and sale practices. [^29] For example, the CCPA secures privacy rights for California consumers, including the right to know about the personal information a business collects about them and how it is used and shared. The GDPR in Europe also requires firms to disclose how they use data, apart from granting consumers more control over data. See Lizzeri (1999) for a similar observability assumption. One interesting special scenario is that in the contract the intermediary commits to a binary data policy (share full information or not), which, however, is not optimal for the intermediary from the second half of part 2 in 3 and hence underlines the importance of a flexible data trade policy.
Continuum of Values.
We assume binary product values. One could expect the continuous distributions of product values would affect consumer surplus. For the value distributions $F_{L}(v),F_{H}(v)\in\Delta[v_{L},v_{H}]$, the producer sets a price between $v_{L}$ and $v_{H}$ rather than $v_{L}$ or $v_{H}$ as the product pricing problem becomes $\max_{p}p[1-F_{H}(p)]\mu_{s}+p[1-F_{L}(p)](1-\mu_{s})$. In the setup of binary values with $\mu_{0}>v_{L}/v_{H}$, only high type consumers may get a positive surplus in the product market. But either type of consumer can get a positive one for continuous product values.
Horizontal-vertical Setup.
We model the data-sourcing market as horizontal, while the data-using market is modeled as vertical. The horizontal part endows consumers with incentives to reveal personal information to match their tastes, while the assumption of the vertical element allows the producer to use consumers’ data against their interests in the data-using market. If both markets are modeled as horizontal, then cutoffs for the intermediary’s optimal data trade would change. If both are modeled as vertical, then consumers may have no incentives to share information in the data-sourcing market.
@@ 7 Conclusion
This paper develops an integrated framework of the intermediary’s data sourcing from consumers and data trade to a third party who uses the data. Consumers benefit from revealing information about their tastes to the intermediary for better-matched services. A third party buys data from the intermediary so that it can extract more surplus from consumers in the product market. We characterize the intermediary optimal mechanisms for data sourcing and sales. Due to the cross-market feature of data, the equilibrium data trade depends on the consumers’ horizontal differentiation level in the data-sourcing service market and the vertical one in the product market.
In the intermediary optimal mechanism, the intermediary sells no information or partial information to preserve profits in the service market when consumers’ mismatch cost from a wrong service is low. The low mismatch cost itself cannot sufficiently incentivize consumers to reveal information. The intermediary has to compensate consumers more (by decreasing service fees) to sell full information. However, the loss in service profit outweighs the net gain from selling full information rather than partial or no information. Hence, the intermediary would prefer to sell partial or no information. In contrast, the intermediary can sell full information to the data buyer when consumers’ mismatch cost in service is high. In this case, consumers’ privacy protection is at the lowest level. Moreover, the provisions of service, product, and data by the optimal mechanism can be efficient under certain market conditions.
Additionally, we introduce the intermediary’s partial bargaining power in the data market. The partial bargaining power does not affect the surpluses of the intermediary, producer, or consumers, although it can decrease the amount of information traded by the intermediary with the third party.
Applying the framework to examine the data market regulations, we find that banning data trade may reduce social welfare because it makes it harder to price discriminate in the product market. Instead, regulating the intermediary’s bargaining power can limit the amount of information traded by the intermediary without hurting social welfare. The finding suggests that the regulation of the intermediary’s power is more efficient than the regulation of the intermediary’s data sharing.
Appendix A Appendix: Omitted Proofs
A.1 Proof of 1